# Infinitary combinatorics in condensed mathematics and strong homology

### v1

@@ 5. Conclusion
This might be loosely thought of as the derived version of the question of whether such products are projective (recall that by Proposition 4.2 the answer is no). Just as for that question, instances of this one admit formulations in functional analytic terms (cf. [23, Appendix to III]). For example:
@@
As indicated, questions like these seem to form the essential content of standing questions falling more purely within the field of algebraic topology:
@@ Acknowledgements
A portion of this work was carried out during its authors’ research visits to the Fields Institute’s spring 2023 Thematic Program on Set Theoretic Methods in Algebra, Dynamics and Geometry; we thank both the institute and the program’s organizers for their hospitality and support. We reiterate our thanks to Dustin Clausen and Peter Scholze for their shaping role in this paper’s results, as well as to Lucas Mann for his generosity in clarifying both the details and significance of several of its arguments, and thank both Catrin Mair and Michael Hrušák for discussions benefiting the paper as well.

### v2

@@ 5. Conclusion
This might be loosely thought of as the derived version of the question of whether such products are projective (recall that by Proposition 4.2 the answer is no). Just as for that question, instances of this one admit formulations in functional analytic terms (cf. [25, Appendix to III]). For example:
@@
See [15] and [6] for more on this question; as shown in the latter work, the question is in fact equivalent to the first part of the next one:
@@ Acknowledgements
A portion of this work was carried out during its authors’ research visits to the Fields Institute’s spring 2023 Thematic Program on Set Theoretic Methods in Algebra, Dynamics and Geometry; we thank both the institute and the program’s organizers for their hospitality and support. We reiterate our thanks to Dustin Clausen and Peter Scholze for their shaping role in this paper’s results, as well as to Lucas Mann for his generosity in clarifying both the details and significance of several of its arguments, and thank both Catrin Mair and Michael Hrušák for discussions benefiting the paper as well. We wish lastly to thank the paper’s referee for so thoroughly alert and prompt a reading of it.